FREQUENCY SHIFTS NEAR THE FIRST ORDER PHASE TRANSITION IN NH_{4}Cl

\underline{H. YURTSEVEN}, Department of Physics, Istanbul Technical University, Maslak, Istanbul, Turkey.

We study here our correlations between the volume changes and the frequency shifts for the q[110] mode of NH_{4}Cl for the pressures of 0, 0.5 and 0.95 kbars (first order phase region). By determining the mode Grüneisen parameter for the q[110] mode, we have predicted the ultrasonic frequencies for this mode using the length-change measurements for the NH_{4}Cl crystals from the literature. Our calculated frequencies are in good agreement with the observed frequencies for the q[110] mode in the first order phase region of NH_{4}Cl. This shows that our method of calculating the frequencies from the observed length-change data is reasonable and that we can then predict the observed behaviour of NH_{4}Cl.